A sum of R1 000 is due in ve months' time from now, at a simple interest rate of 12% per annum. The present value of the sum is

[1] R952,38.

[2] R567,43.

[3] R1 050,00.

[4] R625,00.

[5] none of the above


1
Expert's answer
2021-06-09T16:36:38-0400

"principal\times Rate\times Time+principal=1000"

"principal\times 0.12\times 5/12+principal=1000"


"principal=\frac{1000}{1.05}=952.38"


Answer is option 1 of R 952.38